An air bubble rises from the bottom of a well where the temperature is 25°C to the surface where the temperature is 27°C. Find the percent increase in the volume of the bubble if the depth of the well is 5 m. Atmospheric pressure is 101,528 Pascals.    


1
Expert's answer
2021-10-18T11:01:48-0400

"T_1 = 25 +273 = 298 \;K \\

T_2 = 27 + 273 = 300 \;K \\

p_2 = p_{atm}=101528 \;Pa = 101528 \;N/m^2"

Pressure at bottom due to depth of fluid

"p = ρgh \\

p = 10^3 \times 9.81 \times 5 = 49050 \;N/m^2"

Total pressure at bottom

"p_1 = p_{atm} + p \\

= 101528 +49050 \\

= 150578 \;N/m^2 \\

\frac{p_1V_1}{T_1} = \frac{p_2V_2}{T_2} \\

\frac{150578 \times V_1}{298} = \frac{101528 \times V_2}{300} \\

V_2 = \frac{150578}{101528} \times \frac{300}{298} \times V_1 \\

V_2 = 1.4929V_1 \\

% \; increase \; in \; volume = \frac{V_2-V_1}{V_1} \times 100 \\

= \frac{1.4929V_1-V_1}{V_1} \times 100 \\

= 49.295 \; \%"